OpenStudy (anonymous):

what are the formulas to find: area circumference and to find the area of an irregular shape i'm awarding a medal to the first to answer

OpenStudy (lastdaywork):

Area of a circle = pi*r*r Circumference of a circle = 2*pi*r "... to find the area of an irregular shape..." We use integration :)

OpenStudy (anonymous):

so if the radius is 23 for both of them what is the answer?

OpenStudy (lastdaywork):

Try to plug-in the values and find the answer yourself :)

OpenStudy (anonymous):

I did they both came out as 144.44 the I rounded it to 144.4 is that right?

OpenStudy (anonymous):

or area to be 1661.06

OpenStudy (lastdaywork):

area = 1661.902514 circumference = 144.5132621

OpenStudy (anonymous):

now rounded?

OpenStudy (anonymous):

can you help me with the irregular fiqure

OpenStudy (lastdaywork):

I simply used Excel XD are you going to use the concept of significant figures here??

OpenStudy (lastdaywork):

Upload your irregular figure..

OpenStudy (anonymous):

k

OpenStudy (anonymous):

it wont let me upload but its a half circle underneath a triangle so like this C> and there is a straight line going through the triangle stopping b4 the circle which is 4.8 ft long then the half circle is 3.6ft wide

OpenStudy (anonymous):

is that tooo confusing im sorry lol :(

OpenStudy (lastdaywork):

Take a snapshot and upload the image..Its hard to answer this way..

OpenStudy (anonymous):

r u still there lastdaywork ?

OpenStudy (anonymous):

ihave to find the area

OpenStudy (lastdaywork):

To call people, use @ before their names..like @Hahanopers12321

OpenStudy (anonymous):

ok

OpenStudy (lastdaywork):

Can you find the area of the triangle ??

OpenStudy (anonymous):

isn't it the number x pi?

OpenStudy (anonymous):

@lastdaywork

OpenStudy (anonymous):

like 4.8 x 3.14

OpenStudy (lastdaywork):

Area of triangle = (1/2) * base * height Here base = 3.6 Height = 4.8 & please don't call me when I am already here :P

OpenStudy (anonymous):

what 1/2 ? sorry I am dumb

OpenStudy (anonymous):

so just 4.8 x 3.6

OpenStudy (anonymous):

and thats the area?

OpenStudy (lastdaywork):

Area of triangle = (base * height)/2 Does that makes sense ^^ & dont be sorry, we are all dumb..in a way XD

OpenStudy (lastdaywork):

Total area = area of triangle + area of semi circle

OpenStudy (anonymous):

thanks a lot

OpenStudy (lastdaywork):

:)